### v1

Abstract
Unmanned Aerial Systems (UASs) or drones become more and more commercially available and cheap. The availability of these cheap commercial off the shelf (COTS) UASs in recent years led to potentially more threats to the traditional perimeter defense of civilian and military facilities, critical infrastructures and public events. Moreover, Law Enforcement Agencies (LEAs) find themselves confronted with the novel task of having to police the access to the lower airspace as more and more drones become available. Equally, there has been much emphasis on developing and deploying Counter-UAS systems (UASs) with Detection Tracking and Identification (DTI) solutions. However, the capabilities of these systems are hard to benchmark. Performance claims of these systems are currently not supported by evidence. In addition, no standard test methodologies are available for these DTI systems and different test methodologies make comparison of these systems hard or impossible. We report on the definition, development and verification of an objective-driven test method and corresponding comparative performance evaluation for commercial DTI solutions for C-UASs. The developed methodology is based on end-user scenarios that are operationally relevant. The test methodology is based on a generic DTI system lay-out and is detailed towards detection, tracking and identification, taking into account contextual information and end-user input. The comparative performance evaluation is developed to enable the use of the methodology in a relevant environment, thereby taking into account any potential environmental aspect that might influence DTI system performance. A comprehensive process has been outlined to develop and implement the objective-driven test method and the comparative performance evaluation. This process deals with the design of the testing system where defining the objectives the tests are addressed unambiguously. These objectives are formulated in an objective function that deals with key performance indicators for the system under test. To ensure that tests happen under comparable conditions a controlled testing environment is set up. This controlled environment covers relevant parameters and real-world aspects needed for a test scenario. Finally, such a system has to be improved and validated iteratively, meaning that once completed the test system should be applied to trials where a variety of systems are tested and compared. The work presented in this paper has firstly been verified in a simulation environment where a number relevant scenarios were used and the output the simulation injected into the testing system. Validation of the work in a relevant environment has been done in two operational trials. The operational trial results show that the method allows for performance evaluation at component level (i.e., detection, tracking or identification component) and at system level (combinations of these components and integrated DTI system of system solutions).
keywords:
[…]
1 Introduction
In recent years, remote controlled toys (such as cars or boats and even planes) have matured to become semi-autonomous unmanned vehicles Famili et al. (2024). Advances in virtually all sectors connected to the design, the production and the operation of such vehicles have resulted in commercially available products becoming vastly more capable (e.g., flight time, precision, autonomy, payloads) while at the same time getting cheaper and easier to operate. What used to be considered high-tech equipment just a few years ago Joint Research Centre and Institute for the Protection and Security of the Citizen and Boucher (2014) (e.g., a hovering drone with a reasonable camera providing a stable video feed of a location) is now available for very little money, shipping included, delivery to your doorstep.
In the Ukraine, UAVs have become a consumable resource (as opposed to a prized possession), to be used (quite literally sometimes) like an artillery shell. In this conflict, the majority of drones is currently operated by first-person view (FPV), meaning they are piloted (as opposed to autonomous). With the proliferation of such FPV-vehicles and the growing community of experienced pilots comes the reality that also our civilian airspaces are increasingly subject to such vehicles Tadić et al. (2024). It is just a question of time until unmanned assets, especially Unmanned Aerial Vehicles (UAVs) are also used for nefarious purposes European Commission and Joint Research Centre et al. (2023); Seidaliyeva et al. (2024).
As a result, the industry has reacted by advancing the field of detection, tracking and identification (DTI) systems. These systems, integral to both civilian and defense applications European Commission and Joint Research Centre et al. (2023), range from advanced radio detection and ranging (RADAR) setups to sophisticated software solutions for UAV oversight. Despite their technological advancements and increasing adoption, the evaluation of DTI systems remains a complex challenge in need of objective benchmarks, comprehensive performance metrics, and adaptable testing frameworks. With commercial interests at stake, competition may be fierce and any consultant advising a customer for one and against another system (a) better have their facts straight and (b) be prepared to defend this decision against disgruntled commercial parties.
On top of that, there is the fact that, just as drones, DTI systems can be used in a broad range of circumstances or scenarios Joint Research Centre and Institute for the Protection and Security of the Citizen and Boucher (2014), ranging from e.g., the needs of a hobbyist airport’s traffic controller over a hospital worried about the safety of their emergency helicopter’s airspace to a mayor fearing for the safety of their St. Patrick’s Day parade, not to mention military and defense applications. Within all these different domains a plethora of potential use-cases and scenarios can be imagined, all with their own settings and requirements. For a fair assessment of which system is best suited for a particular job the interested buyer (the user) must be able to include these details into the performance evaluation.
This paper introduces a comprehensive testing methodology tailored to the nuanced requirements of DTI systems. By weaving together a framework that prioritizes objective-driven test methods with the comparative analysis of system performance, our approach aims to offer a versatile and rigorous means of evaluating DTI systems. Through a meticulous integration of key performance indicators Key Performance Indicators (KPIs), weighting of these KPIs, and scenario-based testing, this methodology not only facilitates a nuanced assessment of DTI capabilities but also empowers the testers to tailor the evaluation process to their specific operational needs.
1.1 Motivation
As UASs or drones become more and more commercially available and cheap, Law Enforcement Agencies (LEAs) find themselves confronted with the novel task of having to police the access to the lower airspace. The availability of cheap commercial off the shelf (COTS) UASs in recent years Joint Research Centre and Institute for the Protection and Security of the Citizen and Boucher (2014) has indeed led to potentially more threats European Commission and Joint Research Centre et al. (2023) to the traditional perimeter defense of civilian and military facilities, critical infrastructures and public events Tan et al. (2021). Equally, there has been much emphasis on developing and deploying Counter-UAS systems (C-UASs) with DTI solutions. However, the capabilities of these systems are hard to benchmark.
These C-UASs provide a wide range of capabilities including DTI functionalities. Performance claims of these systems are currently not supported by evidence. No standard test methodologies are available and different test methodologies make comparison of these systems hard or impossible. Each operational environment may require different DTI capabilities or configurations. This may lead to DTI systems performing differently given distinct use case scenarios and environments. The need for the LEAs to have an objective driven test methodology that enables comparative performance evaluation of DTI systems. It is of paramount importance for the LEAs to have a such a test method including performance evaluation at their disposal.
1.2 Scope
This work deals with the definition, development and verification of an objective-driven test method and comparative performance evaluation for commercial DTI solutions for C-UASs. For the development of a test method that addresses relevant end-user objectives, following key elements have been taken into account Mohamoud et al. (2019):
A set of standard end-user defined scenarios representing a wide set of use cases (e.g., prison & airport security, critical infrastructure protection, border security);
Operational end-user needs to mitigate the threat posed by UASs (i.e., drones); and
Functional and performance requirements in close collaboration with the end-users.
Based on these elements, an integral test method including comparative performance evaluation is developed that allows for a fair comparison between different DTI systems. This test method has been verified during three user-scripted demonstration trials. In doing so, the test method fosters a better understanding among LEAs of the capabilities and performance of DTI systems in relevant operational use cases.
1.3 Contributions of this paper
The test method we propose, building upon the operational needs and the functional requirements, enables testing and performance evaluation of DTI systems. Indeed the test method allows for performance evaluation at component level (i.e., detection, tracking or identification component) and at system level (combinations of these components and integrated DTI system of system solutions). Moreover the test method including performance evaluation takes the following main elements into account:
Development of a test method for testing of integrated DTI systems and their sub systems under realistic conditions and using relevant end-user defined scenarios.
A method that enables the end-users to evaluate a DTI system and thereby helps addressing whether the system meets operational needs and requirements from the end-user perspective.
A future-proof methodology which means that the test methodology can be adjusted to test and evaluate future DTI systems.
1.4 Paper Structure
Figure 1 provides a visual overview over the paper. We start by providing the necessary background with regard to objective driven and comparative performance testing in Section 2 and discuss building blocks of DTI systems in Section 3, which also discusses the methodology. This continues into Section 4 where we report on our approach to realize objective-driven, comparative performance evaluation for commercial DTI solutions. Section 5 touches on how we validated our approach and Section 6 closes with a conclusion and some reflections.
2 Objective driven and comparative performance testing
2.1 Introduction
When designing a performance evaluation procedure for a specific class of commercial systems, two fundamental concerns should be addressed: the process should be objective driven and the evaluation should be comparative. Both are necessary to ensure that the so evaluated systems are compared in an efficient and effective manner that is suitable to reveal whether these systems meet relevant operational needs and inform e.g., purchasing decisions towards a specific customer or end-user need.
2.1.1 Objective driven test methods
Objective-driven test methods are structured around clearly defined goals and outcomes that a product, system, or process premisses to achieve. This begins with the identification of the specific objectives the end-user has in mind, examples for this include assessing the functionality of a system, evaluating the system’s performance, verifying system reliability, or stress-testing security aspects of a system. Depending on the focus of the performance evaluation, this should then guide the subsequent design and development of test criteria and procedures as well as the definition of performance metrics. By focusing on the objectives of the evaluation process from the end-user perspective we can ensure that the outcomes align with the system’s intended purpose and the end-user requirements.
An objective driven test method should first identify the key functional requirements of the to be tested (and compared) systems and consider appropriate performance benchmarks for the application. The key functional requirements are derived from relevant operational needs. In order to test different objectives relatively independently test scenarios should be designed. By subjecting a system to these scenarios all requirements and critical features can be evaluated. In addition, the decision to use objective driven performance evaluation facilitates a transparent process that can be understood and agreed upon by developers, stake-holders and end-users.
2.1.2 Comparative performance evaluation
In comparative testing the performances of a number of test subjects or systems are compared, using a standardized set of criteria. It creates, in fact, a rating or an ordering of the tested entities depending on clearly defined operational needs; this allows for the results to be used directly for informed decisions with regard to the choice or purchase of system over the others. Comparative testing can be used to facilitate a transparent, fair and unbiased comparison of the tested entities by subjecting them to identical tests and test situations and by doing so under controlled conditions.
By designing scenarios and test cases that reflect the intended deployment of the system in the tester’s specific use-case the performance comparison does compare systems with a focus on the aspects relevant to the end-user. In many cases, the ultimate decision for a system and against another will be a trade-off, meaning that it is often not the case that a single system outperforms all others in all aspects and conditions. It is very likely that performance in one category comes with either a detrimental impact on another or with an exceedingly high cost. Often one does not need the best system at all cost but simply a system that is good enough. Comparative testing enables the tester to choose from those options which yield the highest return, meaning deliver the desired service while all relevant aspects (e.g., performance, cost) are taken into account.
2.1.3 Objective driven, comparative performance testing
As discussed above, both objective driven testing as well as comparative performance evaluation have significant benefits when comparing systems, software, or processes. While the former focuses on assessing whether specific considerations by the tester are met, the latter compares the performance of multiple systems against a set of tester defined benchmarks. Both methodologies contribute significantly to the making of informed decisions, however, their differences in focus and application underscore the importance of using them carefully to complement each other.
Objective driven testing is particularly relevant in early stages of a project as it forces the tester to define the objectives at the outset. This inherently goal-oriented approach therefore drives clarity and facilitates transparent goals. It is introspective in that it focuses on the system’s capability to meet its intended purpose. Comparative testing on the other hand is entirely concerned with the measurable and quantifiable performance of a system in certain conditions as it will compare different systems in the basis of these metrics. While the former ascertains that a system can indeed perform the services expected from it, the latter compares the degree with which these services are performed across systems.
This is by no means to suggest that these two approaches are mutually exclusive, far from it: both can be used in combination with great effect. For example, when designing a process to facilitate the testing of a class of systems, the design of this process should be objective driven and result in a comparative performance evaluation.
2.2 Important process stages
The process of designing and developing a comparative performance evaluation system can be split in three high level stages: the design of the testing system, the actual use of the system (e.g., by means of an opeational trial) and, ultimately, the validation of the system; cf. Figure 2.
With regard to the design of such a system or process, we already discussed above the need for clearly stated objectives to unambiguously define what the goals of the testing process are. With that in place, we can discuss the metrics we want to use. These metrics must cover all aspects the tester (user) considers relevant for their eventual choice. Examples for such metrics are e.g., system performance and reliability, used hardware, response time of sensors, ease of installation, ease of use or cost (both OPEX as well as CAPEX).
To ensure that tests happen under comparable conditions a controlled testing environment needs to be set up. This controlled environment should have all relevant parameters and real-world aspects needed for a test scenario (of which there are a number) covered. With regard to the testing scenarios, these should be designed to be generic and abstract enough to cover entire classes of groups of real-world scenarios but have to be specific enough to collectively cover the breadth of scenarios, including special cases and outliers. Note that scenarios should be designed to be tunable to specific needs and interests. For example, a scenario may revolve around a train station, in which case the scenario should be an abstract representation of such a building. It should be tunable to a small station during peak traffic as well as to a major station in the middle of the night.
With the above in place, a number of commercially competing products can be subjected to tests and their resulting performances can be compared. Note that preferences for specific metrics over others (maybe reliability is more important than cost-effectiveness) are expressed by the tester and are to be reflected in an objective function which is applied to all partial results for each tested system.
Finally, such a system has to be improved and validated iteratively, meaning that once completed the system should be applied to trials where a variety of systems are tested and compared. Emphasis should be placed in the correctness of the outcome as well as the transparency of the process as well as the easy of use.
2.3 System Design Principles
Underlying the work are some fundamental principles and criteria. These are: comprehensiveness, fairness, and adaptability as well as usability and accessibility.
In order to be broadly applicable, the evaluation procedure must be comprehensive in that it is capable of evaluation a range of performance metrics and can so over a number of diverse scenarios. This should cover the application domain and spectrum of tested systems and represent all major use-cases (in abstract scenarios). The evaluation should be impartial and void of inherent biases (towards specific manufacturers, systems, paradigms or users) and should therefore be able to use any metric and run any scenario when testing a system; metrics and scenarios may not favor any individual system under investigation.
In an area as active and young as that of UASs, which is rapidly evolving, any testing system must be designed such that it can be amended to include changes in the field. Adaptability ensures that the evaluation system remains relevant as novel systems are introduced or when priorities change. This requires that metrics and scenarios can be updated or amended without it affecting the overall architecture of the system.
Finally, the evaluation system has to be available (both in access to the system as well as in ease of use thereof) to the broader community interested in the evaluation of a specific system. If access to the performance comparison is restricted to certain stake-holders then the main goal (that of providing a objective means to fairly assess the performance of systems to a specific use case) cannot be accomplished.
2.4 Defining the objective
An objective function in the context of evaluating complex systems like DTI systems quantitatively represents the user’s priorities and performance aspects they wish to focus on. This objective function’s result is a weighted sum of normalized performance metrics, where each weight corresponds to the user’s emphasis (preference?) for that metric. Users begin by identifying KPIs relevant to their operational needs, such as detection and track accuracy, and identification reliability. Each metric is then assigned a weight based on its importance to the user’s specific application or mission requirements. The evaluation system provides a user-friendly interface for this process, offering templates for common use cases while allowing custom configurations. The objective function is designed to be flexible, accommodating changes in user priorities and the introduction of new KPIs as DTI technologies evolve.
The overall aim is to design an evaluation system that can be tailored by a user to compare commercial systems for a broad set of specific use-cases and scenarios. A user must therefore be able to tune the scenario to represent their use case and must further be able to change the scoring of the system to reflect the user’s priorities and preferences. The latter is captured in the objective function of the performance evaluation, which quantitatively represents the user’s priorities and focuses on those performance aspects the user wishes to focus on.
The objective function commonly uses normalized test scores for specific metrics, weighing them according to the user’s preferences. This function should be configurable for the user, and the system should facilitate this in an easy and intuitively understandable manner.
2.5 Scenario modelling
A number of abstract, but representative scenarios must be modelled. These have to reflect the diverse application domain of the system class to be tested, allowing for specific incarnations of a scenario that covers outlier cases as well as standard situations.
3 A methodology to evaluate DTI solutions for counter-UAS systems
3.1 Operational needs and scenario modelling
Relevant operational needs and requirements form the basis in defining functional requirements the DTI systems should meet. The key underlying operational requirements can be summarized as the need to detect and counter any drone, of any size, exhibiting certain behaviour in different environmental conditions. Scenario modelling for evaluating complex systems like DTI involves the creation of a comprehensive set of environments and conditions under which the systems’ performances are assessed. These scenarios are designed to mimic a broad spectrum of real-world operational conditions, including various climates, terrains UAS malicious and neutral behaviour, and urban density levels. Each scenario is characterized by specific parameters such as weather conditions, the presence of obstacles, and the density of non-target entities (e.g., birds or non-malicious drones) to assess systems’ adaptability and reliability under varied conditions. The process incorporates input from industry experts, operational data, and historical incident reports to ensure relevance and comprehensiveness. This approach ensures that the evaluation system can test DTI systems’ performance in controlled yet realistic and relevant conditions.
The proposed objective driven and comparative evaluation of the DTI systems is built upon a set of generic scenarios. These high-level scenarios can be tailored for the LEA specific needs. The scenarios are grouped into three main domains, specifically:
Sensitive Sites and Critical National Infrastructure (CNI)
Public Spaces Protection and Events
Border Protection (Land and Maritime)
Please see Section 4.3 for some more details on our scenarios.
The operational needs and requirements derived from these scenarios are then translated into a set of functional and performance requirements for DTI systems.
3.2 Functional requirements
Using operational requirements and generic scenarios, a set of functional requirements are derived (these are not addressed in this paper). The operational needs translate into functional requirements for a DTI system that can detect, track, identify (and classify) any drone, of any size, exhibiting certain behaviour in different relevant environmental conditions. Moreover, these functional requirements provide a baseline against which DTI systems tested and evaluated. DTI system functional breakdown, expected behaviour of the system and its components are essential for defining relevant set of performance evaluation metrics, cf. Section 3.4.
3.3 Unmanned aerial systems
Unmanned aerial systems UASs are technologically advanced machines that operate without a human operator physically present on the system. These systems can perform tasks autonomously, relying on pre-programmed instructions and algorithms, or be controlled remotely by a human operator from a distant location Joint Research Centre and Institute for the Protection and Security of the Citizen and Boucher (2014). The key characteristic of unmanned systems is their ability to execute operations with varying degrees of independence from human intervention, which is made possible through a combination of sensors, data processing capabilities, and actuation mechanisms. In the context of developing objective driven and comparative evaluation of DTI systems, UAS also known as drones – may exhibit malicious behaviour and are as such deemed to pose a security threat in the C-UAS domain.
3.4 DTI building blocks
The DTI process is a means to create or increase situational awareness. In particular, awareness with regard to, in our case, (aerial) unmanned systems (drones) Joint Research Centre and Institute for the Protection and Security of the Citizen and Boucher (2014). Therefore, the process can be said to follow three steps / pass through three stages: the initial realization that there is something there (detection), the subsequent observation of that something over a period of time (tracking) which establishes the detected something as an object existing in the world, and, finally, the classification of this object (identification) according to some taxonomy of interest, e.g., friend versus foe or plane versus bird. This classification can also be of the type “is coming towards me” or “the approaching object has malicious intent”.
As described above and as shown in Figure 3, these three steps are sequential.
3.4.1 Functional decomposition of DTI systems
The functional requirements mainly address the ability to detect an drone at a given range, translate the detections into a track over a period of time, classify the drone based on the track information and identify the drone and its potential threat Broek et al. (2019). As a matter of fact, a DTI can be decomposed into three main functional components, shown in Figure 3:
- 1.
2.
Tracking functionality (tracking of objects – also known as object assessment – builds upon the output of the detection functionality and can be done using a single sensor or a combination of various sensors)
3.
Identification functionality (includes classification of UAS and its potential threat)
3.4.2 Detection
Detection Famili et al. (2024) is the first step in a DTI process. In it, the system identifies sufficient evidence in its sensory input to conclude that there is an object present or at least that there is some phenomena in the real world that has caused the sensors to measure data (as opposed to noise) Seidaliyeva et al. (2024). This is a single observation and in itself not very useful. Depending on the scenario, there may be many detected objects in the sensed environment at any given time.
The technologies (sensors) required to perform detection can come from a variety of fields Seidaliyeva et al. (2024): common detection sensors include RADAR, light detection and ranging (LIDAR), sound navigation and ranging (SONAR), infra-red (IF)- or radio frequency (RF)-sensors, or optical cameras. The identification of evidence for the existence of an object is challenging because of the multitude of signals that can be received by a sensor in an environment. This noise has to be filtered and the remaining signal processed, the sensitivity of the sensors (as well as their range) have to be tuned to the scenario (and the object one wishes to detect) to avoid false positives while detecting as many true positives as possible (see Section 3.5.1).
3.4.3 Tracking
If an object has been detected once, the next step is to detect it (i.e., the same object) again, which can be rather challenging when there are a number of objects in the environment. In a way, tracking is connecting the dots, i.e., the recognition of a specific object so that one can label the dots representing its individual detections. Together with detections from previous moment in time these individual detections can turn into a track. The track of a detected object represents the object’s path through the environment. It might exhibit some behaviour and may be indicative of the object’s nature and / or intent.
In addition, a precise track for an object possibly allows for the projection of the object’s path. If, for example, the object’s mass, velocity and air resistance are known then one might be able to calculate the resulting trajectory of the object.
The number of detections in a track may be relevant. For example, a few dots will allow the drawing of a ragged line which would be much smoother if more dots (in-between the drawn ones) were available. If there are enough dots then the displaying only the last one with a high frequency might give the illusion of the dot moving on the screen. However, effective tracking through(quasi) continuous observations requires maintaining detection as the object moves, and to do so in a manner that enables the system to infer that the new detection and the previous one are in fact from the same object. Consequently, tracking essentially requires detection (or sensors that facilitate detection) as well as software to process the received data and to infer which detection belongs to which object.
3.4.4 Identification
The identification process deals with the classification of an object based on its track and certain characteristics or attributes. This is the logical next step after detecting (becoming aware of) an object and tracking (observing) it. Classification can refer to learning more about the nature and identity of the object but may also extend to inferring the object’s intent. With regard to DTI systems, this can mean identifying the nature (e.g., bird, drone, plane), the type (e.g., UAV, Unmanned Ground Vehicle (UGV)) or the sub-type (e.g., quad-copter, fixed-wing) of an object, but can also extend to guessing the object’s reason for being in the area (e.g., hobbyist flying their drone, criminals attempting to deliver drugs to a prison or spies collecting information about the perimeter) or inferring the object’s intent (such as e.g., hostile or benign).
| Aspect | Explanation |
| Type Classification | Determining the type or sub-type of a detected object. Examples are distinguishing between different types of drones, birds, or other airborne objects. This might extend to recognizing characteristics that offer insight into a drone’s payload (capability) or purpose (intent). |
| Friend or Foe Determination | In the domain of defense, safety and security offering a high-level classification with regard to an object being friendly, hostile, or neutral (or unknown) is a crucial first step in the decision process behind planning a response. |
| Behavioral Analysis | Identification can involve an analysis of the observed behaviour or a matching of observed patterns to known modi operandi. Examples are low flying drones which may want to avoid detection or devices loitering in a specific area which may indicate a surveillance activity. |
| Feature Recognition | Attributing an object to some entity can be greatly improved when specific and possibly identifying features of the detected object can be observed. Examples for such features are shape, size or color but can also be signal emissions or used communication protocols. Advanced DTI systems might use machine learning algorithms to improve feature recognition and identification accuracy over time. |
| Registration and Comparison | Finally, for high value or high interest objects there may be a database with identifying features that enables a system to match an observed object uniquely to a catalogue of known objects. While maybe not the perfect example for a DTI system, it is a known fact that navy vessels have a unique signature in their noise profile as well as their electromagnetic field. A system could e.g., uniquely identify any individual US aircraft carrier by its signatures. |
3.5 Performance metrics
Performance metrics are designed to quantitatively assess a system’s effectiveness, efficiency, reliability, and usability Kouhestani et al. (2017). Table 2 shows example performance metrics for evaluating complex DTI systems.
| Metric focus | Example metrics |
| Effectiveness | Detection range; accuracy; the system’s ability to identify and track targets under various conditions. |
| Efficiency | The system’s resource utilization, such as energy consumption and processing time. |
| Reliability | The system’s operational consistency over time and under different scenarios, including its resilience to failures or environmental challenges. |
| Usability | The system’s ease of use, including setup time, required training, and user interface intuitiveness. |
These metrics should be standardized to allow a fair comparison across different systems but must also be customizable to reflect the specific priorities and needs of the tester. In the next section (3.5.1) we first introduce standard performance metrics before discussing the metrics of our DTI system in specific in Sections 3.5.2 (Detection), 3.5.3 (Tracking) and 3.5.4 (Identification).
3.5.1 Standard performance metrics
Generically speaking, and independent of the application, the a few metrics are commonly used and are intuitively understandable. These can be used anywhere where there is (a) a prediction / classification, (b) a known ground truth (meaning that we do have all relevant data with regard to the actual state of affairs) and (c) more than a few results are available (meaning: where a representative number of tests have been performed). To lay the foundations for this we define true/false positives/negatives:
| Predicted | Predicted | ||||||
| Negative | Positive | not a cat | is a cat | ||||
| Actual | Negative | True Negative | False Positive | not a cat | True Negative | False Positive | |
| Positive | False Negative | True Positive | is a cat | False Negative | True Positive | ||
For example, we may have 1000 pictures and know precisely which of these pictures shows a cat. We can then ask a toddler to identify all the pictures of a cat and use the responses to create 4 piles of pictures (see below and Figure 3). Using the number of pictures in each pile, we can now the assess the child’s precision and recall performance.
True Negative: images correctly identified as not having a cat in them.
False Positive: images incorrectly identified as having a cat in them.
False Negative: images incorrectly identified as not having a cat in them.
True Positive: images correctly identified as having a cat in them.
Precision and recall
Both performance measures look at the number of correctly identified cat pictures (yes, we are sticking with the example), i.e., the true positives. Precision compares this number to the number of all pictures identified as having a kitten in them (both, true and false positives), while recall compares it to the number of pictures that actually have a cat in them (true positives + false negatives).
F1 score
The above provided definitions for precision and recall can be easily exploited: the former incentivise being extremely cautious (ideally selecting a single picture that then does feature a cat, to get a perfect score of ) while the latter is even easier: simply classify everything as a kitten picture and, since the incorrect predictions are not used in the formula, this guarantees a score of . Both measures have their merit but the silly examples above make it clear that stand-alone these values only look good on paper.
What is needed is a measure that uses both, precision and recall in order to force a trade-off. The community has long since decided that this shall be captured in the so-called F1 score, defined as follows:
A few notes here: both, precision as well as recall are percentages, meaning both will always be between 1 and zero. This means that in the above formula the dividend will always be smaller than 1, and will also always be smaller than the divisor. Therefore, the F1 score will also always be a value between 1 and zero. In addition, the formula rewards similar scores for precision and over one significantly exceeding the other.
False and Missed Alarms
Given the application we may want to include to more (very similar) metrics: precision and recall measure how well a system is performing, and for applications like machine learning this is indeed what we want to know. For our performance evaluation process we want the user to be able to decide what their priorities are and where their preferences lie. We therefor also want to include measures for negative performance, such as Missed Alarm Rate (MAR) and False Alarm Rate (FAR) (cf. Table 6):
3.5.2 Detection
With regard to the detection aspect of commercial DTI systems we considered 3 metrics, namely the accuracy of the location of the detected drone (both in 2D as well as in 3D), the range ratio for detection and the precision of the detection. See Table 4.
| Metric name | Description |
| Location accuracy (2D/3D) | The location accuracy of a detection representing a true object is defined as the distance between the detection and the true object. The metric is undefined for detections which do not represent a true object. |
| Range ratio | The relative minimum and maximum detection range of a true object is defined as the minimum and maximum distance of the detections representing the object from the DTI system normalized for the minimum and maximum range of the true object within the Area of Interest (AoI) from the DTI system. |
| Precision | The precision of detections is defined as the fraction of all detections which represent a true object, cf. Equation 1. |
3.5.3 Tracking
To evaluate the performance of a DTI system with regard to the tracking of drones we decided upon 7 metrics: Track completeness, continuity, ambiguity and spuriousness as well as the accuracy of the track with regard to velocity and position. Finally, we included the longest track segment as a metric as well. See Table 5 for more details.
| Metric name | Description |
| Track completeness | The track completeness of a true object is defined as the fraction of time in which the object is represented by at least one track. |
| Track continuity | The track continuity of a true object is defined as the total number of tracks representing the object. The metric is undefined if the true object has no tracks representing the object. |
| Track ambiguity | The track ambiguity of a true object is defined as the time-weighted average of the number of tracks representing the object during the time the object has at least one track representing the object. The metric is undefined if the true object has no tracks representing the object. |
| Track spuriousness | The track spuriousness is defined as the time-weighted average of the number of tracks not representing a true object at that time. |
| Track velocity accuracy | The track velocity accuracy of a true object is defined as the Root Mean Square (RMS) velocity difference between the tracks representing the object and the true object. The metric is undefined when no track represents the true object. |
| Track positional accuracy | The track positional accuracy of a true object is defined as the RMS distance between the tracks representing the object and the true object. The metric is undefined when no track represents the true object. |
| Longest track segment | The longest track segment of a true object is defined as the largest fraction of time in which the object was represented by the same track while being in the AoI. |
3.5.4 Identification
For the identification capabilities of a DTI system we agreed on 4 separate metrics: F1 to capture precision and recall performance, FAR, MAR and Probability of Detection (PoD); these are briefly explained in Table 6.
| Metric name | Description |
| F1 | The F1-score is the harmonic mean of precision and recall, cf. Equation 3. |
| False Alarm Rate (FAR) | The false alarm rate is defined as the fraction of falsely given alarms out of the total number of alarms, cf. Equation 5. |
| Missed Alarm Rate (MAR) | The missed alarm rate is defined as the fraction of alarms the DTI did not emit out of the total number of alarms it should have emitted, cf. Equation 4. |
| Probability of Detection (PoD) | The probability of detection is defined as the number of times in which the DTI system emits the alarm rightfully (true positives), divided by the total number of alarms it should have emitted; cf. Equation 2. |
4 An objective-driven, comparative performance evaluation for DTI solutions
4.1 Performance evaluation pipeline
The performance evaluation is based on:
Use of relevant end-user defined test scenarios that contain, for instance, contextual information and objects (e.g., border, CNI) that are to be secured.
A set of evaluated performance metrics that, starting from the results of a test or a collection of tests, provide a score to the DTI system under test and its components.
A rating that is updated every time a new test iteration has been executed.
The performance evaluation pipeline consists of two main parts, namely, Metric level, and Component level evaluation.
4.1.1 The metric level
The metric level starts with tests being executed based on the scenarios. The measurements emanating from these tests are captured and translated into a set of metrics. At the metric level, the performance of a specific metric (e.g., detection range, track continuity) is evaluated. These metrics are then interpreted and normalized into a score. In order to know how to score the desired metric, a ‘scoring context’ can be added. This scoring context provides information about what is operationally desired for a given metric in a given context. In the end, the normalized score between 0 and 1 is calculated, where 1 is the best possible score and 0 the worst score. See Figure 4 for an illustration.
4.1.2 The component level
At the component level, normalized scores calculated at the metric level are used. The component test level gathers all scores belonging to a specific functionality (e.g., detection, tracking, identification). These scores are optionally weighted based on operational end-user prioritization and then aggregated. The result is an aggregated score for that specific functionality. A rating is created based on the scores of single tests, on the scores per capability and KPIs based on operational needs. The rating is updated every time a DTI system is tested. This is shown in Figure 5.
The aggregated scores emanating from various component level results in the case different DTI systems and subsystems are evaluated can be combined. In doing so, the full functionality of an integrated DTI system, given the DTI is comprised of a set of (sub)systems can be evaluated and compared. This combination of aggregated DTI performance scores from this DTI components allows for a comparative performance evaluation that enables evaluation of integrated DTI system of systems.
4.1.3 System Architecture
The evaluation system’s architecture is designed to be modular, scalable, and adaptable, ensuring it can accommodate a wide range of complex commercial DTI systems and evolving evaluation needs. At its core, the architecture consists of three main components: the User / End-user Input, Scenario Database and the Evaluation Engine.
The User Input provides an accessible platform for users to define their objective functions, select or customize evaluation scenarios, and view results. It is designed with usability in mind, catering to users with varying levels of technical expertise. The Evaluation Engine is the system’s computational heart, processing objective functions, running DTI systems through selected scenarios, carrying out evaluation of DTI systems though performance metrics. It utilizes a modular design, allowing new performance metrics and evaluation algorithms to be integrated as they are developed. The Scenario Database houses a comprehensive collection of pre-defined scenarios and allows for the storage of custom scenarios created by users. This component supports the system’s adaptability, enabling it to grow and evolve in response to new insights, technologies, and user requirements. Together, these components form a robust system capable of evaluating complex commercial DTI systems in a fair, comprehensive, and user-driven manner.
4.2 Defining the Objective Function
The definition of operationally relevant objective function is of vital importance for the comparative evaluation of DTI systems. From the end-user perspective a DTI system evaluation should be based on its ability to detect, track and identify potential malicious drones in a given AoI. This is also of paramount importance for the subsequent mitigation measures including countering measures which are out of the scope of this work. The guiding aspects in the context of the comparative performance evaluation are testing and evaluation of integrated DTI systems and their sub systems under realistic conditions and using relevant end-user scenarios and KPIs. To this end, depending on the operational requirements, specific use case scenarios and the available DTI functionality, a set performance metrics and end-user determined weighting factors are defined. The objective is defined such that its result is a weighted sum of normalized performance metrics where each weight corresponds to the user’s emphasis. The end-user prioritization through the use of weighting factors can be applied to a single metric, combined metrics at component level or metrics at integrated system level. The evaluation system including the objective function provides the user with the ability to formulate how well a DTI system should cover specific operational requirements.
4.3 Scenario Modelling
The proposed objective driven and comparative evaluation of the DTI systems is built upon a set of end-user defined operational needs and a set of scenarios. The scenarios are grouped into three categories:
1.
Protection of Sensitive Sites and Critical National Infrastructure (e.g., prison, airport, nuclear plant, government building)
2.
Protection of Public Spaces and Events (e.g., a stadium, outdoor events)
3.
Border Protection (e.g., land border and maritime border)
The end-user defined scenarios cover different operational needs and come with varying contextual and environmental conditions. The common denominator for these scenarios are:
The definition of specific objects to be secured, including area of interest.
The definition of type threats (drones) that is to be expected (e.g., above ground level information, speed, size of drone and its type. Both commercial and custom-made drones are considered)
The expected behaviour the threat (drones) such as direction the threat is coming from, the entry into the area of interest
Contextual and environmental aspects (e.g., urban, suburban and rural, day and night, weather conditions and terrain)
The user-defined operational requirements (e.g., at which range a drone must be detected and an alarm generated. This depends on the specific scenario parameters)
5 Validation of methodology
A validation approach for the objective driven and comparative performance evaluation is presented here. The proposed approach is two-fold:
1.
A simulation-based validation approach to authenticate the developed test methodology and corresponding comparative performance evaluation in a simulation environment that facilitates sensitivity tests.
2.
A trial-based validation approach to confirm the objective driven test methodology and comparative performance evaluation in an operational environment with end-user defined relevant scenarios.
5.1 Test framework
A basic concept of a test framework has been defined building upon the object driven test methodology and comparative performance evaluation presented in XX. This framework enables the validation of the proposed performance evaluation. The test framework depicted in XX contains the following main components:
Test environment dealing with scenario’s including environmental aspects (e.g., trees, birds, buildings) and objects of interest (drones).
System under evaluation representing the DTI system that is to be tested and evaluated. For the simulation-based validation, a basic low fidelity model of a DTI system has been created with detection, tracking and identification strategies (i.e., alarm generating) capability. The purpose of the low fidelity model is to allow for verifying the comparative performance evaluation prior to its use in the trial-based validation.
Test suite covering the test methodology and corresponding performance evaluation using the defined metrics.
5.2 Test environment
The test environment can be virtual (a simulation environment) or a physical test terrain where the DTI system under test can be deployed for testing purposes (i.e., trial-based validation). An underlying requirement for the evaluation of the DTI system under test is the inclusion of representative environment which includes objects of interest (e.g., neutral traffic and malicious drones) to be detected and classified. The concept also facilitates the storing of test data (e.g., ground truth data of objects in the environment, output of DTI systems like detections, tracks, and identification) and logging of relevant settings (e.g., position of DTI systems).
The test environment enables evaluation of the DTI system in its actual environment. The interaction between the DTI system and the main relevant aspects of the environment is shown in Figure 7. The DTI interaction with its environment contains the following components:
DTI system representing the detection, tracking and identification system that is under test in a relevant environment.
Various relevant factors from the environment that can influence the DTI system performance (e.g., weather, buildings, birds, aircraft, terrain and trees). Moreover, the environment can be polluted by electromagnetic (EM) emissions (e.g., Wireless LAN, 4G).
Drones can be present in the environment that exhibit certain behaviours (e.g., friendly or neutral, malicious). The drones can also generate EM emissions. Drones are detected based on their emissions and on the signatures they have when illuminated.
Location of interest depicting the area to be observed by the DTI system which is linked to a specific scenario (e.g., border scenario, critical infrastructure).
5.3 Test Suite
5.3.1 Simulation Test Environment
A test suite tailored for the simulation-based validation has been developed. This test suite given in FIGURE XXX enables the execution of a number of standardized tests and generation of stimuli. After the execution of the tests, evaluation of the results of the system under test is carried out. The main components of the test suite are:
Predefined series of tests with a range of scenarios that are executed under controlled conditions in various vignettes to derive metrics generated from results of system under test (DTI output). These metrics are used to evaluate the system in a standardized and repeatable fashion. A test-coordinator is responsible for selecting which test is executed, preparing the test environment, starting the test itself and retrieving results.
Performance evaluation process that – based on the output of the system under test during the standardized tests – comes up with a scoring of the system of interest on all tested levels.
From the predefined test sets, the scenarios to run and the stimuli to be created are known. Based on this information, multiple runs of the same test can be carried out in the test suite. From each of these runs, output of the DTI model is generated and stored for the evaluation process. In this process, first the metrics are calculated and subsequently evaluated taking into account end-users’ operational context. The evaluated metrics can be aggregated to come to an overall score for the DTI in that specific test scenario.
The test steps that cover all the necessary elements for the validation method are given in Figure 8.
These test steps are:
Test configuration: the testing starts by loading a pre-defined test configuration.
Load environment: A configuration file that defines which scenarios are run and how many iterations per scenario are required.
Test execution: during an iteration, a test is executed. Execution of the test in the test environment means generating stimuli (neutral and malicious drones), obstacles (trees, buildings) and providing this information to the available DTI models.
Store data: The output of the DTI models is stored for analysis.
Performance analysis: Carry out comparative performance evaluation based on the DTI output. In case the iterations of a scenario are done, a new scenario is loaded, and new tests are performed, until all tests specified in the test configuration are handled.
5.3.2 Operational Trial Data Processing
The main objective of the trial-based validation is to authenticate the developed objective-driven test methodology and corresponding comparative performance evaluation using real world operational data. To make this possible, a data processing and analysis is defined for the trial-based validation methodology. For the trial-based validation, operational trials are organized and executed using relevant scenarios (relevant environment, drones exhibiting both neutral and malicious behaviour) and DTI systems. For the processing analysing of data, and subsequent evaluation of the DTI systems, the following steps can be identified (reference in Figure 10, below):
Collection of DTI output (e.g., detections, tracks, identification output) and ground truth data from the flying objects (e.g., drones) from the trial data.
Parsing of the collected data (i.e., DTI output and ground truth).
Pre-processing of the data (e.g., time window filtering, transformation of coordinate system and selection of AoI).
Association of detections and tracks to the ground truth data.
Calculation and normalization of performance metrics.
Inclusion of end-user context (e.g., weighting of metrics).
Generation of an aggregated score per DTI.
6 Discussion
6.1 Discussion
We report here an objective test methodology and a comparative performance evaluation for DTI. The developed methodology is based on end-user scenarios that are operationally relevant. The test methodology is based on a generic DTI system lay-out and is detailed towards detection, tracking and identification, taking into account contextual information and end-user input. The comparative performance evaluation is developed to enable the use of the methodology in a relevant environment, thereby taking into account any potential environmental aspect that might influence DTI system performance.
The developed objective-driven test methodology and comparative performance evaluation has been verified using a two-fold approach:
A simulation-based approach to validate the developed test methodology and corresponding performance evaluation. This is done in a simulation environment that facilitates carrying out sensitivity tests.
A trial-based validation approach to confirm the objective driven test methodology and comparative performance evaluation in an operational environment with end-user defined relevant scenarios.
For the simulation-based validation, several runs were executed to generate stimuli data as input for the test methodology and the corresponding comparative performance evaluation. Using the low fidelity DTI model, detections (object at a certain location at a given time), track (multiple detections fused into one track over time), and alarms (in case an alarm was given by the DTI) had been generated. Several metrics were generated from the output of the DTI and the stored ground truth data that contains the actual positions of the simulated drones. The following metrics were used in the test runs:
Range ratio
Location accuracy 2D
Track completeness
F1 score
The test runs were repeated for three different DTI model strategies and different characteristics (such as field of view and track association capabilities). As the characteristics of the DTI model strategies are known, validation of the performance evaluation results is possible.
For the trial-based validation, DTI output data (detection, tracking and identification capabilities) from two operational trials have been utilized. A number of relevant end-user defined operational scenarios have been carried out (an example of the drone flight paths – the ground truth - is given in Figures 11 and 12). Note that the DTI output obtained during the trials is, per the agreement with the participating companies, confidential and therefore not included in this report. Various companies participated in the trials, fielding DTI systems covering RADAR, EO††margin: abbrev /IF, RF capabilities and combinations thereof, which were subjected to the presented test methodology in these trials. The complete processing and evaluation pipeline depicted in Figure 11 has been covered. The paramount purpose of using the operational data is to validate the performance evaluation pipeline.
Using the operational data showed together with the performance evaluation pipeline, the following steps have been successfully carried out:
Correct parsing of the collected data (both DTI output such detections, tracks and identifications and drone ground truth).
Pre-processing of the data (e.g., time window filtering, transformation of coordinate system and selection of AoI).
Association of detections and tracks to the ground truth data (DTI track to ground-truth track association).
Calculation of performance metrics.
Inclusion of end-user context (e.g., weighting of metrics for certain metrics).
Generation of an score per DTI at metric level and aggregated score at component level and system level.
More than 20 DTI systems’ output has been collected during the operational trials and processed (and example of relevant ground truth data is given in Figures 11 and 12). Various performance metrics at different levels (metric, component, and combinations of these) have been calculated and evaluated. The trial-based validation results show the objective-driven test methodology and the comparative performance evaluation successfully process operational data, calculate performance metrics and generate (aggregated) scores for DTI (sub)systems while taking into account the end-user input.
Conceptualization, A.M. and R.vH. methodology, A.M. H.H. and R.vH. software, J.vdP. B.M. M.vdH. and A.vK. validation, A.M. J.vdP. and M.vdH. formal analysis, investigation, A.M. J.vdP. H.H. R.vH. B.M. M.vdH. and A.vK. resources, A.M. data curation, J.vdP. M.vdH. and A.vK. writing—original draft preparation, A.M. and H.H. writing—review and editing, A.M. J.vdP. H.H. R.vH. B.M. M.vdH. and A.vK. visualization, J.vdP. H.H. M.vdH. and A.vK. supervision, A.M. R.vH. and B.M. project administration, A.M. and B.M. funding acquisition, A.M.
All authors have read and agreed to the published version of the manuscript.
This work has been carried out in project COURAGEOUS. This project has received funding from the European Union’s Internal Security Fund Police (ISFP) under Grant Agreement 101034655
“Not applicable”. No drones were harmed during the trials.
“Not applicable”.
As per the agreement with the participating companies, the performance data for specific systems under specific test conditions is considered confidential and cannot be made available publicly by the authors.
Acknowledgements.
[…]
Acronyms
- AoI
- Area of Interest
- C-UAS
- Counter-Unmanned Aerial System (UAS) system
- CNI
- Critical National Infrastructure
- COTS
- commercial off the shelf
- DTI
- detection, tracking and identification
- EM
- electromagnetic
- FAR
- False Alarm Rate
- FPV
- first-person view
- THUAS
- The Hague University of Applied Sciences
- IF
- infra-red
- KPI
- Key Performance Indicatior
- LEA
- Law Enforcement Agency
- LIDAR
- light detection and ranging
- MAR
- Missed Alarm Rate
- PoD
- Probability of Detection
- RADAR
- radio detection and ranging
- RF
- radio frequency
- RMS
- Root Mean Square
- SONAR
- sound navigation and ranging
- TNO
- Netherlands Organisation for Applied Scientific Research
- UAS
- Unmanned Aerial System
- UAV
- Unmanned Aerial Vehicle
- UGV
- Unmanned Ground Vehicle
[…]